# Accelerated Gradient Methods with Biased Gradient Estimates: Risk Sensitivity, High-Probability Guarantees, and Large Deviation Bounds

### v1

@@ 4.1. Numerical experiments comparing risk, convergence rate, and parameter choice
In this section, we investigate the relationship among risk-sensitive cost, parameter selection, and convergence rate in common parameterizations of GMMs from a numerical perspective. See table˜1 for a list of common parameters which covers HB, NAG and GD methods. The last column of this table provides the linear convergence rate $\rho\in(0,1)$ associated with the stated parameters as a function of the condition number $\kappa:=L/\mu$ when there is no noise.If $w_{k+1}=0$ for every $k$; then $\rho^{2}\in[0,1)$ is such that $f(x_{k})-f(x_{*})\leq\bar{c}_{k}\rho^{2k}[f(x_{0})-f(x_{*})]$ where $\bar{c}_{k}$ has at most polynomial growth, see e.g. [58]. Here, we abuse the notation and use $\rho$ to denote both the spectral radius of a matrix and the convergence rate, similar to the literature [94, 57].
@@
We focus on a simple strongly convex quadratic problem of the form (4.1) with $d=2,L=3,\mu=1,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g=[0,~0]^{\top}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h}=0,Q=\begin{bmatrix}\mu&0\\ 0&L\end{bmatrix}$. We set $\sigma^{2}=2$, and we conduct the experiments for various values of $\theta>0$ chosen appropriately so that $\sqrt{\theta}H_{\infty}<\sqrt{d}=\sqrt{2}$ –for otherwise $R(\theta)$ would be infinite by theorem˜4.1.
@@
In fig.˜1, we study risk as $\theta$ varies for various common GMM parameterizations. Theorem 4.1 provides an exact formula for the risk index $R(\theta)$ when $\theta\in\left(0,\frac{d}{H_{\infty}^{2}}\right)$ and shows that the risk index is infinite if $\theta\geq\bar{\theta}:=\frac{d}{H_{\infty}^{2}}$.
@@
In LABEL:fig:_HB-risk-rate, we consider the HB algorithm on a quadratic function with $\mu=1,L=3,d=2,\sigma^{2}=2,\theta=1.2$ and plot the logarithm of risk-sensitive index (LABEL:fig:_HB-risk-stepsize) and the convergence rate $\rho$ (LABEL:fig:_HB-rate-stepsize) as a function of the step size $\alpha$ and momentum parameter $\beta$ –see [58, Lemma 3.1] for a precise convergence rate characterization of HB for strongly convex quadratics. The black dot indicates the standard HB parameters that yield the fastest convergence rate, while the blue cross represents the parameters of the RS-HB method, which achieve the best known rate while minimizing the $H_{\infty}$ norm [57]. The shaded gray region denotes parameter values for which the risk becomes infinite. As shown in LABEL:fig:_HB-risk-rate, the fastest convergence rate in the noiseless case (black dot) comes at the cost of higher risk, whereas the RS-HB configuration (blue cross), offers improved robustness with a lower risk sensitivity and $H_{\infty}$ norm. Overall, we observe that faster convergence rates (lighter colors in LABEL:fig:_HB-rate-stepsize) are typically accompanied by higher risk (darker colors in LABEL:fig:_HB-risk-stepsize).
In LABEL:fig:_NAG-risk-rate, we observe a phenomenon similar to the HB case for the NAG algorithm. The parameter setting labeled NAG-fastest (black dot), which achieves the fastest convergence rate in the absence of noise, incurs a higher risk compared to NAG-pop (blue cross), a commonly used parameterization. While NAG-fastest offers faster convergence, it comes at the cost of increased risk. As in GD and HB, the higher risk values are generally associated with faster convergence rates also for NAG.
@@
As demonstrated so far, a faster rate, i.e., smaller $\rho\in(0,1)$, comes often at the expense of an increased risk sensitivity index. To better understand this relationship, fixing the risk parameter $\theta=0.2$, we examine the Pareto-optimal boundary in Figure˜3 for the same simple quadratic problem we used for fig.˜1, i.e., $d=2,L=3,\mu=1$, and $\sigma^{2}=2$. In this setting, a point $(\rho,R)$ in the plane of convergence rate $\rho\in(0,1)$ and risk sensitivity index $R\geq 0$ is said to lie on the Pareto-optimal boundary if there is no other achievable pair $(\rho^{\prime},R^{\prime})$ with $\rho^{\prime}<\rho$ and $R^{\prime}\leq R$, or $\rho^{\prime}\leq\rho$ and $R^{\prime}<R.$ Equivalently, one cannot improve both the rate and the risk simultaneously beyond a Pareto-optimal point. In fig.˜3, each smooth curve corresponds to one algorithm (GD, HB, NAG), parametrized by its step-size $\alpha$ (and momentum parameters $\beta,\nu$ where applicable). As we slide $\alpha$ (and $\beta,\nu$) within the stability region, we trace out the set of all achievable $\bigl{(}\rho(\alpha,\beta,\nu),\,R(\alpha,\beta,\nu)\bigr{)}$. The Pareto frontier is the upper-left envelope of these curves: it marks the parameters that has the lowest risk $R$ for a given achievable rate $\rho$ for a method. This frontier is numerically estimated via a grid search over the parameter space using the convergence rate formula from [58] and the explicit risk characterization provided in theorem˜4.1. A $3000\times 3000$ grid is used for HB and NAG over the interval $(0,\frac{2}{L})\times(0,2)$ for $(\alpha,\beta)$. For HB we set $\nu=0$, while for NAG we set $\nu=\beta$. A grid of size 5000 is used for GD over the interval $(0,\alpha^{*})$, where $\alpha^{*}\coloneqq\frac{2}{L+\mu}=\max\{\frac{1}{L},\frac{2}{L+\mu},\frac{2}{L+\sqrt{L\mu}}\}$, for which GD achieves the fastest rate.
@@
The dashed vertical lines on the right of Figure˜3 marks the standard rate $1-\frac{1}{\kappa}$ achieved by gradient descent with common step size $\alpha=1/L$ while the other line on the left marks the accelerated rate $1-\frac{1}{\sqrt{\kappa}}$ with the square root dependency to $\kappa$; indeed, $1-\frac{1}{\kappa}>1-\frac{1}{\sqrt{\kappa}}$ here as $\kappa>1$. Dots label classical parameter choices within the parameter space of each algorithm –see Table˜1 for the definitions of these parameter choices. We observe that GD achieves the most conservative frontier: its Pareto-optimal curve lies above those of the momentum methods HB and NAG shown here. Although GD-fastest attains the best rate achievable by GD, it remains slower and riskier than NAG-pop and NAG-fastest. For moderate rates ($\rho\approx 0.5$$1.0$), the Pareto-optimal curve of HB shows that HB can achieve lower risk than GD. At very fast rates ($\rho\approx 0.25$$0.275$) which GD cannot achieve, HB’s accelerated dynamics come with increased risk sensitivity. NAG’s popular tuning, NAG-pop, lies slightly below HB for low-to-moderate rates, achieving a small reduction in risk for the same $\rho$. NAG-fastest is faster than NAG-pop as expected, but admits a slightly higher risk than what HB can achieve for the same rate. On the other hand, TMM sits clearly above NAG and HB, indicating that while it is able to achieve a faster rate than GD, its risk sensitivity index is more than its counterparts which are able to achieve the same rate (NAG-pop has exactly the same rate) in the quadratic setting. Finally, if the step size, $\alpha$, is carefully selected, one can achieve the same rate as GD-pop with NAG-beta-opt while simultaneously minimizing risk sensitivity among all competing popular parameterizations –indeed, for Figure˜3 we selected $\alpha=\frac{\mu}{L^{2}}$ for NAG-beta-opt to match GD-pop’s convergence rate.
@@ 4.2. Large deviations and the risk-sensitive index
It is well-known that averaging iterates can help with optimization performance in inexact gradient settings, see e.g. [57, 98, 99, 24, 111] and the references therein. If we consider the ergodic averages of the iterates
@@
where $E^{o}$ and $\bar{E}$ denotes the interior and closure of the set $E$. If only the right-most inequality in (4.16) is satisfied, then we say that $\Delta_{K}$ satisfies an LDP upper bound. The function $I(\cdot)$ is called the rate function. A rate function is called a good rate function if it admits compact level sets; this allows the infimum over $I(s)$ in (4.16) to be attained at some point in the closure of $E$, i.e., $\bar{E}$. One difficulty in characterizing a good rate function arises from the fact that the random variables $\{f(x_{k})-f_{*}\}_{k\geq 0}$ are not i.i.d., even when the noise sequence $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{w_{k+1}\}_{k\geq 0}$} is i.i.d. Consequently, standard large deviations results such as Cramér’s Theorem do not apply. In the following result, we characterize the rate function by verifying the conditions of the Gärtner–Ellis Theorem [64], which accommodates non-i.i.d. sequences similar to $\{f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\}_{k\geq 0}$. This characterization further leverages the risk-sensitive index formula, established in the proof of theorem˜4.1, expressed as an integral over the unit circle.
@@
$R(\theta)=\begin{cases}-\frac{\sigma^{2}}{2\pi\theta}\displaystyle\int_{-\pi}^{\pi}\log\det\left(\mathsf{I}_{d}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\theta}{d}}G(e^{i\omega})G^{*}(e^{i\omega})\right)d\omega{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}&\mbox{if}\quad\theta\in(-\infty,0)\mbox{ or }\theta\in(0,\frac{d}{H_{\infty}^{2}}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};}\\ \frac{\sigma^{2}}{2\pi d}\displaystyle\int_{-\pi}^{\pi}\operatorname{trace}\left(G(e^{i\omega})G^{*}(e^{i\omega})\right)d\omega,&\mbox{if}\quad\theta=0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};}\\ +\infty,&\mbox{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}otherwise}},\end{cases}$ (4.18)
with $G(z)\coloneqq T(z\mathsf{I}_{2d}-A_{Q})^{-1}B$ for $z\in\mathbb{C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\setminus}\mbox{\text{Spec}}(A_{Q})$ where $\mbox{\text{Spec}}(A_{Q})$ denotes the set of eigenvalues of $A_{Q}$, $G^{*}(e^{i\omega})$ denotes the Hermitian (conjugate) transpose of $G(e^{i\omega})$, $B$ and $T$ are defined in (3.4) and (4.7), respectively. Furthermore, for any $t>0$,
@@
In the setting of proposition˜4.3, the formula (4.18) is equivalent to
$R(\theta)=\begin{cases}-\frac{\sigma^{2}}{2\pi\theta}\displaystyle\sum_{j=1}^{d}\displaystyle\int_{-\pi}^{\pi}\log\left(1-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\theta}{d}}h_{\omega}(\lambda_{i})\right)d\omega{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}&\mbox{if}\quad\theta\in(-\infty,0)\mbox{ or }\theta\in(0,\frac{d}{H_{\infty}^{2}}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};}\\ \frac{\sigma^{2}}{2\pi d}\displaystyle\sum_{j=1}^{d}\displaystyle\int_{-\pi}^{\pi}h_{\omega}(\lambda_{i})d\omega{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}&\mbox{if}\quad\theta=0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};}\\ +\infty,&\mbox{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}otherwise}},\end{cases}$ (4.20)
where $\{\lambda_{i}\}_{i=1}^{d}$ are the eigenvalues of $Q$,
@@
with $\tilde{b}_{\lambda}$, and $\tilde{c}_{\lambda}$ defined by (3.17).
@@
In LABEL:fig:rate-fn-quadratics, we illustrate proposition˜4.3 for a quadratic objective as in (4.1) in dimension $d=2$ with parameters $g=[0,0]^{\top}$, $h=0$, and $Q=\begin{bmatrix}\mu&0\\ 0&L\end{bmatrix}$ where $\mu=1$, $L=3$. The figure displays the rate function $I(t)$ corresponding to various methods and parameter choices within the GMM framework, as listed in table˜1, using the dual representation (4.17). This entails computing $R(\theta)$, for which we evaluate the integral in (4.18) numerically. For $\theta>0$, one may alternatively compute $R(\theta)$ based on the alternative representation  (4.11). Since $I(t)$ is a good rate function, it has compact level sets, a property that is evident in the plots. Given that the gradient noise is i.i.d. Gaussian, the iterates follow a Gaussian distribution, and the average suboptimality $S_{K}/(K+1)$ exhibits exponential tails for all $K$, including in the limit as $K\to\infty$. This behavior results in the linear growth of $I(t)$ in the large-$t$ regime. Notably, in this regime, HB-fastest, the fastest in terms of convergence rate (see table˜1), has the smallest values of $I(t)$, which indicates a slower decay of probabilities for large deviations. As seen from formula (4.17), the rate function is, up to scaling, the convex conjugate (Fenchel–Legendre transform) of the risk-sensitive index (see the proof of proposition˜4.3 for further details). Consequently, a similar trade-off between convergence rate and risk arises: faster convergence entails a slower decay of the probabilities of rare events for sufficiently large $t$. Due to the parabolic shape of $I(t)$, its minimum is attained at the most probable value of $t$, i.e., smaller or larger values being relatively less likely.

### v2

@@ 4.1. Numerical experiments comparing risk, convergence rate, and parameter choice
In this section, we investigate the relationship among risk-sensitive cost, parameter selection, and convergence rate in common parameterizations of GMMs from a numerical perspective. See table˜1 for a list of common parameters which covers HB, NAG and GD methods. The last column of this table provides the linear convergence rate $\rho\in(0,1)$ associated with the stated parameters as a function of the condition number $\kappa:=L/\mu$ when there is no noise. [^2] If $w_{k+1}=0$ for every $k$; then $\rho^{2}\in[0,1)$ is such that $f(x_{k})-f(x_{*})\leq\bar{c}_{k}\rho^{2k}[f(x_{0})-f(x_{*})]$ where $\bar{c}_{k}$ has at most polynomial growth, see e.g., [58]. Here, we abuse the notation and use $\rho$ to denote both the spectral radius of a matrix and the convergence rate, similar to the literature [95, 57].
@@
We focus on a simple strongly convex quadratic problem of the form (4.1) with $d=2,L=3,\mu=1,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g=[0,~0]^{\top}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}h}=0,Q=\begin{bmatrix}\mu&0\\ 0&L\end{bmatrix}$. We set $\sigma^{2}=2$, and we conduct the experiments for various values of $\theta>0$ chosen appropriately so that $\sqrt{\theta}H_{\infty}<\sqrt{d}=\sqrt{2}$ –otherwise $R(\theta)$ would be infinite by theorem˜4.1.
@@
In fig.˜1, we plot $R(\theta)$ for several GMM parameterizations, computed analytically with the DARE-based formula given in theorem˜4.1. This formula is valid for $\theta\in\left(0,\tfrac{d}{H_{\infty}^{2}}\right)$, and $R(\theta)$ diverges when $\theta\geq\bar{\theta}:=\tfrac{d}{H_{\infty}^{2}}$; hence, we truncated each curve in fig.˜1 around $\bar{\theta}$.
@@
(a) (a) $H_{\infty}$ as a function of $\alpha$ based on (4.13) in the same problem setting; (b) The risk-sensitive index as a function of GD step-size $\alpha$ for $\theta\approx\{4/3,~4\}$, where the objective is a quadratic with $d=2,L=3,\mu=1$, and $\sigma^{2}=2$.
@@
(b) (a) Logarithm of the risk-sensitive index $R(\theta)$ as a function of the parameters $\alpha,\beta$ of HB with $\theta=1.2$, $\sigma^{2}=2$$\nu=0$. (b) Convergence rate $\rho$ as a function of $\alpha,\beta$. The objective is a quadratic with $d=2,L=3,\mu=1$.
In fig.˜3(b), we consider the HB algorithm on a quadratic function with $\mu=1,L=3,d=2,\sigma^{2}=2,\theta=1.2$ and plot the logarithm of risk-sensitive index (LABEL:fig:_HB-risk-stepsize) and the convergence rate $\rho$ (LABEL:fig:_HB-rate-stepsize) as a function of the step size $\alpha$ and momentum parameter $\beta$ –see [58, Lemma 3.1] for a precise convergence rate characterization of HB for strongly convex quadratics. The black dot indicates the standard HB parameters that yield the fastest convergence rate, while the blue cross represents the parameters of the RS-HB method, which achieve the best known rate while minimizing the $H_{\infty}$-norm [57]. The shaded gray region denotes parameter values for which the risk becomes infinite. As shown in fig.˜3(b), the fastest convergence rate in the noiseless case (black dot) comes at the cost of higher risk, whereas the RS-HB configuration (blue cross), offers improved robustness with a lower risk sensitivity and $H_{\infty}$-norm. Overall, we observe that faster convergence rates (lighter colors in LABEL:fig:_HB-rate-stepsize) are typically accompanied by higher risk (darker colors in LABEL:fig:_HB-risk-stepsize).
In fig.˜3(c), we observe a phenomenon similar to the HB case for the NAG algorithm. The parameter setting labeled NAG-fastest (black dot), which achieves the fastest convergence rate in the absence of noise, incurs a higher risk compared to NAG-pop (blue cross), a commonly used parameterization. While NAG-fastest offers faster convergence, it comes at the cost of increased risk. As in GD and HB, the higher risk values are generally associated with faster convergence rates also for NAG.
(c) (a) Logarithm of the risk-sensitive index $R(\theta)$ as a function of the parameters $\alpha,\beta$ of NAG with $\theta=3.7$, $\sigma^{2}=2$$\nu=\beta$. (b) Convergence rate $\rho$ as a function of $\alpha,\beta$, where $\beta_{*}(\alpha)=\frac{1-\sqrt{\alpha\mu}}{1+\sqrt{\alpha\mu}}$ for $\alpha\in[0,\frac{1}{L}]$. The objective is a quadratic with $d=2,L=3,\mu=1$.
@@
As demonstrated so far, a faster rate, i.e., smaller $\rho\in(0,1)$, comes often at the expense of an increased risk sensitivity index. To better understand this relationship, fixing the risk parameter $\theta=0.2$, we examine the Pareto-optimal boundary in Figure˜3 for the same simple quadratic problem we used for fig.˜1, i.e., $d=2,L=3,\mu=1$, and $\sigma^{2}=2$. In this setting, a point $(\rho,R)$ in the plane of convergence rate $\rho\in(0,1)$ and risk sensitivity index $R\geq 0$ is said to lie on the Pareto‐optimal boundary if there is no other achievable pair $(\rho^{\prime},R^{\prime})$ with $\rho^{\prime}<\rho$ and $R^{\prime}\leq R$, or $\rho^{\prime}\leq\rho$ and $R^{\prime}<R.$ Equivalently, one cannot improve both the rate and the risk simultaneously beyond a Pareto‐optimal point. In fig.˜3, each smooth curve corresponds to one algorithm (GD, HB, NAG), parametrized by its step‐size $\alpha$ (and momentum parameters $\beta,\nu$ where applicable). As we slide $\alpha$ (and $\beta,\nu$) within the stability region, we trace out the set of all achievable $\bigl(\rho(\alpha,\beta,\nu),\,R(\alpha,\beta,\nu)\bigr)$. The Pareto frontier is the upper‐left envelope of these curves: it marks the parameters that has the lowest risk $R$ for a given achievable rate $\rho$ for a method. This frontier is numerically estimated via a grid search over the parameter space using the convergence rate formula from [58] and the explicit risk characterization provided in theorem˜4.1.444A $3000\times 3000$ grid is used for HB and NAG over the interval $(0,\frac{2}{L})\times(0,2)$ for $(\alpha,\beta)$. For HB we set $\nu=0$, while for NAG we set $\nu=\beta$. A grid of size 5000 is used for GD over the interval $(0,\alpha^{*})$, where $\alpha^{*}\coloneqq\frac{2}{L+\mu}=\max\{\frac{1}{L},\frac{2}{L+\mu},\frac{2}{L+\sqrt{L\mu}}\}$, for which GD achieves the fastest rate.
@@
The dashed vertical lines on the right of Figure˜3 marks the standard rate $1-\frac{1}{\kappa}$ achieved by gradient descent with common step size $\alpha=1/L$ while the other line on the left marks the accelerated rate $1-\frac{1}{\sqrt{\kappa}}$ with the square root dependency to $\kappa$; indeed, $1-\frac{1}{\kappa}>1-\frac{1}{\sqrt{\kappa}}$ here as $\kappa>1$. Dots label classical parameter choices within the parameter space of each algorithm –see Table˜1 for the definitions of these parameter choices. We observe that GD achieves the most conservative frontier: its Pareto-optimal curve lies above those of the momentum methods HB and NAG shown here. Although GD-fastest attains the best rate achievable by GD, it remains slower and riskier than NAG-pop and NAG-fastest. For moderate rates ($\rho\approx 0.5$$1.0$), the Pareto-optimal curve of HB shows that HB can achieve lower risk than GD. At very fast rates ($\rho\approx 0.25$$0.275$) which GD cannot achieve, HB’s accelerated dynamics come with increased risk sensitivity. NAG’s popular tuning, NAG-pop, lies slightly below HB for low‐to‐moderate rates, achieving a small reduction in risk for the same $\rho$. NAG-fastest is faster than NAG-pop as expected, but admits a slightly higher risk than what HB can achieve for the same rate. On the other hand, TMM sits clearly above NAG and HB, indicating that while it is able to achieve a faster rate than GD, its risk sensitivity index is more than its counterparts which are able to achieve the same rate (NAG-pop has exactly the same rate) in the quadratic setting. Finally, if the step size, $\alpha$, is carefully selected, one can achieve the same rate as GD-pop with NAG-beta-opt while simultaneously minimizing risk sensitivity among all competing popular parameterizations –indeed, for Figure˜3 we selected $\alpha=\frac{\mu}{L^{2}}$ for NAG-beta-opt to match GD-pop’s convergence rate.
@@ 4.2. Large deviations and the risk-sensitive index
It is well-known that averaging iterates can help with optimization performance in inexact gradient settings, see e.g., [57, 99, 100, 24, 114] and the references therein. If we consider the ergodic averages of the iterates
@@
where $E^{o}$ and $\bar{E}$ denotes the interior and closure of the set $E$. If only the (upper-bound) right-most inequality in (4.16) holds (either for all Borel sets $E$ or for certain classes of sets), then we say that $\{\Delta_{K}\}$ satisfies an LDP upper bound. The function $I(\cdot)$ is called the rate function. A rate function is called a good rate function if it admits compact level sets; this allows the infimum over $I(s)$ in (4.16) to be attained at some point in the closure of $E$, i.e., $\bar{E}$. One difficulty in characterizing a good rate function arises from the fact that the random variables $\{f(x_{k})-f_{*}\}_{k\geq 0}$ are not i.i.d., even when the noise sequence $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{w_{k+1}\}_{k\geq 0}$} is i.i.d. Consequently, standard large deviations results such as Cramér’s Theorem do not apply. In the following result, we characterize the rate function by verifying the conditions of the Gärtner–Ellis Theorem [65], which accommodates non-i.i.d. sequences similar to $\{f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\}_{k\geq 0}$. This characterization further leverages the risk-sensitive index formula, established in the proof of theorem˜4.1, expressed as an integral over the unit circle.
@@
$R(\theta)=\begin{cases}-\dfrac{\sigma^{2}}{2\pi\,\theta}\displaystyle\int_{-\pi}^{\pi}\log\det\!\Bigl(\mathsf{I}_{d}-\dfrac{\theta}{d}\,G\!\left(e^{i\omega}\right)G\!\left(e^{i\omega}\right)^{*}\Bigr)\,d\omega,&\text{if }-\infty<\theta<\frac{d}{H_{\infty}^{2}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};}\\[8.61108pt] +\infty,&\text{if }\theta\geq\frac{d}{H_{\infty}^{2}},\end{cases}$ (4.18)
with the convention that the value at $\theta=0$ is interpreted by continuity as the limit
$R(0)$ $=$ $\lim_{\theta\to 0}-\dfrac{\sigma^{2}}{2\pi\,\theta}\displaystyle\int_{-\pi}^{\pi}\log\det\!\Bigl(\mathsf{I}_{d}-\dfrac{\theta}{d}\,G\!\left(e^{i\omega}\right)G\!\left(e^{i\omega}\right)^{*}\Bigr)\,d\omega$ (4.19)
$=$ $\frac{\sigma^{2}}{2\pi d}\int_{-\pi}^{\pi}\operatorname{trace}\!\Bigl(G\!\left(e^{i\omega}\right)G\!\left(e^{i\omega}\right)^{*}\Bigr)\,d\omega.$ (4.20)
Here, $G(z)\coloneqq T(z\mathsf{I}_{2d}-A_{Q})^{-1}B\in\mathbb{R}^{d\times d}$ for $z\in\mathbb{C}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\setminus}\mbox{\text{Spec}}(A_{Q})$ where $\mbox{\text{Spec}}(A_{Q})$ denotes the set of eigenvalues of $A_{Q}$, $G\!\left(e^{i\omega}\right)^{*}$ denotes the Hermitian (conjugate) transpose of $G(e^{i\omega})$, $B$ and $T$ are defined in (3.4) and (4.7), respectively. Furthermore, for any $t>0$,
@@
Remark 4.5.
Since $R(\theta)<+\infty$ if and only if $\theta H_{\infty}^{2}<d$ (cf. (4.18)), the supremum in (4.17) may equivalently be taken over all $\theta\in\mathbb{R}$, i.e., $I(s)=\sup_{\theta\in\mathbb{R}}\frac{\theta}{2\sigma^{2}}\bigl(s-R(\theta)\bigr).$ With the change of variable $\lambda=\theta/(2\sigma^{2})$ and $\tilde{\Lambda}(\lambda)=\lambda\,R(2\sigma^{2}\lambda)$, we obtain $I(s)=\sup_{\lambda\in\mathbb{R}}\{\lambda s-\tilde{\Lambda}(\lambda)\},$ i.e., $I$ is the Fenchel–Legendre (convex) conjugate of $\tilde{\Lambda}$; hence, $\tilde{\Lambda}$ is simply a scaled version of the risk-sensitive index. See section˜C.1 (proof of proposition˜4.4) for details.
@@
In the setting of proposition˜4.4, the formula (4.18) is equivalent to
$R(\theta)=\begin{cases}-\frac{\sigma^{2}}{2\pi\theta}\displaystyle\sum_{i=1}^{d}\displaystyle\int_{-\pi}^{\pi}\log\left(1-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\theta}{d}}h_{\omega}(\lambda_{i})\right)d\omega{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}&\mbox{if}\quad-\infty<\theta<\frac{d}{H_{\infty}^{2}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};}\\ +\infty,&\theta\geq\frac{d}{H_{\infty}^{2}},\end{cases}$ (4.22)
with the convention that the value at $\theta=0$ is interpreted by continuity as the limit
$R(0)=\lim_{\theta\to 0}-\frac{\sigma^{2}}{2\pi\theta}\displaystyle\sum_{i=1}^{d}\displaystyle\int_{-\pi}^{\pi}\log\left(1-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\theta}{d}}h_{\omega}(\lambda_{i})\right)d\omega=\frac{\sigma^{2}}{2\pi d}\displaystyle\sum_{i=1}^{d}\displaystyle\int_{-\pi}^{\pi}h_{\omega}(\lambda_{i})d\omega.$
Here, $\{\lambda_{i}\}_{i=1}^{d}$ are the eigenvalues of $Q$,
@@
with $\tilde{b}_{\lambda}=\alpha\lambda(1+\nu)-(1+\beta)$ and ${\tilde{c}_{\lambda}}=\beta-\alpha\lambda\nu$ as in (3.17).
@@
(a) (a) The rate function, $I(t)$, for common GMM parameterizations as a function of $t$, a given suboptimality threshold. (b) The rate function, $I(t)$, for NAG-beta-opt varying step-size $\alpha$ and varying suboptimality threshold, $t$. The objective is a quadratic with $d=2,L=3,\mu=1,\sigma^{2}=2$.
In LABEL:fig:rate-fn-quadratics, we illustrate proposition˜4.4 for a quadratic objective as in (4.1) in dimension $d=2$ with parameters $g=[0,0]^{\top}$, $h=0$, and $Q=\begin{bmatrix}\mu&0\\ 0&L\end{bmatrix}$ where $\mu=1$, $L=3$. The figure displays the rate function $I(t)$ corresponding to various methods and parameter choices within the GMM framework, as listed in table˜1, using the dual representation (4.17). This entails computing $R(\theta)$, for which we evaluate the integral in (4.18) numerically. For $\theta>0$, one may alternatively compute $R(\theta)$ based on the alternative representation  (4.11). Since $I(t)$ is a good rate function, it has compact level sets, a property that is evident in the plots. Given that the gradient noise is i.i.d. Gaussian, the iterates follow a Gaussian distribution, and the average suboptimality $S_{K}/(K+1)$ exhibits exponential tails for all $K$, including in the limit as $K\to\infty$. This behavior results in the linear growth of $I(t)$ in the large-$t$ regime. Notably, in this regime, HB-fastest, the fastest in terms of convergence rate (see table˜1), has the smallest values of $I(t)$, which indicates a slower decay of probabilities for large deviations. As seen from formula (4.17), the rate function is, up to scaling, the convex conjugate (Fenchel–Legendre transform) of the risk-sensitive index (see remark˜4.5). Consequently, a similar trade-off between convergence rate and risk arises: faster convergence entails a slower decay of the probabilities of rare events for sufficiently large $t$. Due to the parabolic shape of $I(t)$, its minimum is attained at the most probable value of $t$, i.e., smaller or larger values being relatively less likely.